OpenStudy (anonymous):

The square root of 12 times the square root of 18 ((:

OpenStudy (anonymous):

12=2x2x2 and 18=3x3x2 Now solve.:)

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{12}\sqrt{18}\] \[\Large \sqrt{12*18}\] \[\Large \sqrt{216}\] Now simplify

jimthompson5910 (jim_thompson5910):

You can use the idea that 216 = 36*6

OpenStudy (anonymous):

Okay,,, I got 14.69693846

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{216}\] \[\Large \sqrt{36*6}\] Break that up

OpenStudy (anonymous):

@Champs Help a lady out? (:

OpenStudy (anonymous):

36x6 is 216.

jimthompson5910 (jim_thompson5910):

yes, now break up \[\Large \sqrt{36*6}\] and tell me what you get

OpenStudy (anonymous):

Okay bro, you're really starting to confuse me.

OpenStudy (anonymous):

14.6969385

OpenStudy (anonymous):